# Skein modules and character varieties of Seifert manifolds

### v1

@@ 1. Introduction
$\mathcal{X}(M)=\mathrm{Hom}(\pi_{1}(M),\mathrm{SL}_{2}(\mathbb{C}))//\mathrm{SL}_{2}(\mathbb{C}),$
@@
Our results in [DKS23] imply if $X(M)$ is infinite, then $\mathcal{S}(M,\mathbb{Q}[A^{\pm 1}])$ is non-tame. On the other hand, by the Culler-Shalen theory [CS83], $M$ contains an essential surface. Hence, Conjecture 1.1 holds for 3-manifolds with infinite $X(M)$.
@@
Combining Theorem 1.2 with an earlier result of the authors [DKS23, Theorem 1.1], leads to the following corollary proved in Section 4:
@@ 1.1. Outline of contents
The paper is organized as follows: In Section 2 we discuss relations between different, related, properties of skein modules. In Section 3 we prove Theorem 1.3. Section 4 is devoted to the proof of Theorem 1.2. First, using our earlier work [DKS23, Theorem 1.1] and Theorem 1.3, we reduce the proof of Theorem 1.2 to showing that the skein module $\mathcal{S}(M,\mathbb{Z}[A^{\pm 1}])$ is finitely generated for any Seifert fibered space over $S^{2}$ with at most three exceptional fibers and with a non-zero Euler number. We prove the latter statement by skein-theoretic techniques. In Section 5, we compute $|X(M)|$ for all non-Haken Seifert fibered 3-manifolds and, in particular, we prove Theorem 1.5. In the case where $p_{1},p_{2},p_{3}$ are weakly coprime we also compute a basis of $\mathcal{S}(M)$.
@@ 1.2. Acknowledgement
During the course of this work the first author was partially supported by the projects ”NAQI-34T” (ANR-23-ERCS-0008) and by the project “CLICQ” of the Région Bourgogne Franche Comté. The second author was partially supported by the NSF grants DMS-2004155 and DMS-2304033. The third author was partially supported by the Simons Foundation grant 957582.
@@ 2. Relations between different properties of skein modules
    $\mathcal{S}(M)$ has no $(A+1)$- nor $(A-1)$-torsion over $R=\mathbb{Z}[A^{\pm 1}]$.
@@
The implication $(3)\Rightarrow(6)$ follows from the main theorem of [DKS23]. Indeed, if $\mathcal{S}(M)$ is tame, then $\dim_{Q(A)}(\mathcal{S}(M,\mathbb{Q}(A)))\geq|X(M)|,$ therefore $X(M)$ is finite.
For reduced $\mathcal{X}(M)$, the implication $(3)\Rightarrow(5)$ is the last part of [DKS23, Theorem 3.1].
Finally, if $\mathcal{S}(M,\mathbb{Z}[A^{\pm 1}])$ is generated by $|X(M)|$ elements, then it is isomorphic to $\mathbb{Z}[A^{\pm 1}]^{|X(M)|}/I$ where $I$ is some $\mathbb{Z}[A^{\pm 1}]$-submodule of $\mathbb{Z}[A^{\pm 1}]^{|X(M)|}.$ If $I$ is not the trivial submodule, then we would have that $S(M,\mathbb{Q}(A))=\left(\mathbb{Z}[A^{\pm 1}]^{|X(M)|}/I\right)\underset{\mathbb{Z}[A^{\pm 1}]}{\otimes}\mathbb{Q}(A)$ would have dimension $<|X(M)|$ over $\mathbb{Q}(A),$ which contradicts [DKS23, Theorem 3.1]. Hence, we must have $I=0,$ and $(7)\Rightarrow(1).$
@@ 3. Haken Seifert fibered manifolds
We denote by $M(B;\ \frac{q_{1}}{p_{1}}\ldots\frac{q_{n}}{p_{n}})$ the Seifert fibered $3$-manifold with the fiber space being the $2$-orbifold $B$ with exceptional fiber invariants $(q_{1},p_{1})\ldots(q_{n},p_{n})$ in the notation of [JN83].
@@
The representation $\rho$ now satisfies all relations in the presentation 1, and furthermore, maps each $c_{1}\ldots c_{i}$ for $22\leq i\leq n-$ 2 to an element of angle $\varphi_{i}.$
We note that Lemma 3.3 actually implies that the dimension of $X(M)$ is at least $n-3$ for $M$ a Seifert manifold over $S^{2}$ with $n\geq 4$ exceptional fibers. We will however not need this fact, but just that $X(M)$ is infinite. The proof of Proposition 3.1 now reduces to the other case, manifolds which fiber over $\mathbb{R}\mathbb{P}^{2}$ with at least $2$ exceptional fibers:
@@
Since $\frac{q_{1}}{p_{1}}$ and $\frac{q_{2}}{p_{2}}$ are not integers, $\sin\left(\frac{q_{1}}{p_{1}}\pi\right)\sin\left(\frac{q_{2}}{p_{2}}\pi\right)\neq 0.$ Consequently, for different choices of $v_{1},v_{2}$, the trace $\mathrm{Tr}(\rho(c_{1}c_{2}))$ can take any value in the interval $[\alpha-\beta,\alpha+\beta]$, for
@@
Proof of implication :absent:$\Leftarrow$:⇐ : Since $X(S^{2}\times S^{1})$ is infinite, we can assume that $M$ is Haken. Let $(M,\pi,B)$ be a Seifert fibered space structure on $M$ where $B$ is the orbifold of the fibration and $\pi:M\longrightarrow B$ is the canonical projection. Recall that $B$ is a closed surface that may be orientable or non-orientable. Since $\pi$ induces a surjection of $\pi_{1}(M)\to\pi_{1}(B)$, if $B\neq S^{2},\mathbb{R}\mathbb{P}^{2}$, then $H_{1}(M)$ projects onto $\mathbb{Z}$, implying an infinite $X(M)$ in this case.
@@ 4. Non-Haken Seifert fibered manifolds
(c) $\Rightarrow$ (a): By [DKS23, Theorem 1.1], we have $|X(M)|\leq\dim_{\mathbb{Q}(A)}\mathcal{S}(M)$ and since by [GJS23], the dimension $\dim_{\mathbb{Q}(A)}\mathcal{S}(M)$ is finite, we conclude that $X(M)$ is finite. Now by Theorem 1.3, $M$ has to be non-Haken, implying it is either reducible or it contains an incompressible surface. ∎
@@ 4.1. The torus skein algebra and the Frohman-Gelca basis
Since multicurves on $T$ without contractible components form a basis of $\mathcal{S}(T,\mathbb{Z}[A^{\pm 1}])$, the skeins $(p,q)_{T}$ for ±$\pm(p,q)\in\mathbb{Z}^{2}/_{\{\pm 1\}}$ form a basis of $\mathcal{S}(T,\mathbb{Z}[A^{\pm 1}]).$
@@ 4.3. The skein module $\mathcal{S}(S_{0,3}\times S^{1},\mathbb{Z}[A^{\pm 1}]))$
Note that the remaining vectors, $v_{i}$, for $i=3,...,10$, have their first two components $(k_{1},l_{1}).$ Hence, going from $v_{1}$ to any of them decreases $c_{1}/p_{1}$ by $(p_{1}+q_{1})/p_{1}=1+\varepsilon_{1}.$ Therefore, by (7), it is enough to show that going from $v_{1}$ to one of these vectors increases $c_{2}/p_{2}+c_{3}/p_{3}$ by at most $\max(|\varepsilon_{2}-1|+|\varepsilon_{3}|,|\varepsilon_{2}|+|\varepsilon_{3}+1|)$.
@@
Now we complete the proof of Theorem 4.1 as a corollary of Propositions 4.3 and 4.4.
@@
Hence, the above inequality follows from the fact that $|\varepsilon_{3}+1|$ is either $-\varepsilon_{3}-1$ or is less than $1$. Recall that $\mathcal{S}(M)$ is generated by the elements $\bar{L}_{v}$ for $v\in\mathbb{Z}^{6}.$ By Propositions 4.3 and 4.4 (which holds by the above inequality), we see that $\mathcal{S}(M)$ is generated by the elements $\bar{L}_{v}$ satisfying
@@
Combining Theorem 1.2 and an earlier result of the authors [DKS23, Theorem 1.1], leads to Corollary 1.4 which we restate here:
@@
It is a direct consequence of Theorem 1.2, combined with an earlier result of the authors [DKS23, Theorem 1.1]. Although Theorem 1.2 implies tameness for irreducible $3$-manifolds only, the above statement does hold the reducible Seifert $3$-manifolds, $S^{2}\times S^{1}$ and $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$. In the first case, $\dim_{\mathbb{Q}(A)}\mathcal{S}(S^{2}\times S^{1})=1$ by [HP95] and in the latter,
@@ 5. Character varieties of non-Haken Seifert manifolds
Let $R(M):=\mathrm{Hom}(\pi_{1}(M),\mathrm{SL}_{2}(\mathbb{C}))$. A character $\chi$ of $M$ is called abelian if it is the trace of a diagonal representation $\rho\in R(M)$ and central if $\rho$ takes values in the center {±$\{\pm I\}$ of $SL(2,\mathbb{C})$. A character is called irreducible if it is the trace of an irreducible $\rho\in R(M)$.
@@ 5.1. Abelian characters
where $\lambda\in\mathrm{Hom}(\pi_{1}(M),\mathbb{C}^{*})\simeq H_{1}(M,\mathbb{Z}).$ Two such representations $\rho_{\lambda}$ and $\rho_{\lambda^{\prime}}$ are conjugated if and only if $\lambda=(\lambda^{\prime})^{\pm 1}$ (as maps $\pi_{1}(M)\rightarrow\mathbb{C}^{*}$). Since one has $\lambda=\lambda^{-1}$ if and only if $\lambda$ takes values in {±$\{\pm 1\},$ that is, $\lambda\in H_{1}(M,\mathbb{Z}/2\mathbb{Z}),$ the statement follows.
@@
For the second formula, note that for each pair $(p_{i},q_{i})$ at most one of the components is even. So the matrix of above presentation is non-zero mod $2$ and the dimension of $H_{1}(M,\mathbb{Z}/2\mathbb{Z})$ over $\mathbb{Z}/2\mathbb{Z}$ is at most $2$ The dimension is non-zero if and only if the determinant of this matrix is 00 mod $2$ For the dimension to be $2$ one needs all of the $2\times 2$ minors: $p_{1}q_{2},\ q_{1}p_{2},\ p_{1}p_{2},\ -q_{1}p_{3}-p_{1}q_{3},\ -q_{1}p_{3},\ -p_{1}p_{3},\ -q_{2}p_{3},\ -q_{2}p_{3}-p_{2}q_{3},\ p_{2}p_{3}$ to vanish in $\mathbb{Z}/2\mathbb{Z}$. The later happens if and only if $p_{1},p_{2}$ and $p_{3}$ are even. ∎
@@
where $m$ is the number of $\varphi\in\mathrm{Hom}(\pi_{1}(M),\mathbb{C}^{*})$ with $\varphi(h)=\pm 1,$ and $m_{i}$ is the number of $\varphi\in\mathrm{Hom}(\pi_{1}(M),\mathbb{C}^{*})$ such that $\varphi(h)=\varphi(c_{i})=\pm 1.$ (Note that if two of $c_{1},c_{2},c_{3}$ are mapped to ±$\pm 1,$ so is the third one). By Lemma 5.1, $m=|H_{1}(M)/(2h)|$. Eliminating $c_{3}$ from the presentation of $H_{1}(M)/(2h)$ that with respect to the ordered set of generators $\{h,c_{1},c_{2}\}$ the presentation matrix is
@@
If $H_{1}(M,\mathbb{Z})$ is $2$-torsion, then $x_{M}=0$ because all abelian representations have values in {±$\{\pm I\}$ in this case.
@@ 5.2. Irreducible characters
Each $\chi\in X_{2}(M)$ is either irreducible or exceptional abelian. In either case, $\rho(h)=I$ and $\chi$ is determined by its values on $F_{2}=\langle c_{1},c_{2},c_{3}|c_{1}c_{2}c_{3}=1\rangle$. Since $F_{2}$ is free of rank two, $\chi$ is determined by $(\chi(c_{1}),\chi(c_{2}),\chi(c_{3}))\in\mathbb{C}^{3}$. Thus $T_{+}$ is $1$-$1$. Furthermore, each such value in $\mathbb{C}^{3}$ determines a character of $F_{2}$. In particular, each value in $C_{p_{1}}^{+}\times C_{p_{2}}^{+}\times C_{p_{3}}^{+}$ corresponds to a character of $\langle c_{1},c_{2},c_{3}|c_{1}^{p_{1}}=c_{2}^{p_{2}}=c_{3}^{p_{3}}=c_{1}c_{2}c_{3}=1\rangle$. Thus $T_{+}$ is onto. The same argument shows that $T_{-}$ is a bijection.

### v2

@@ 1. Introduction
$\mathcal{X}(M)=\mathrm{Hom}(\pi_{1}(M),\mathrm{SL}_{2}(\mathbb{C}))/\hskip-3.61371pt/\mathrm{SL}_{2}(\mathbb{C}),$
@@
Our results in [DKS25] imply if $X(M)$ is infinite (for example, then $\mathcal{S}(M,\mathbb{Q}[A^{\pm 1}])$ is non-tame. On the other hand, by the Culler-Shalen theory [CS83], $M$ contains an essential surface. Hence, Conjecture 1.1 holds for 3-manifolds with infinite $X(M)$.
@@
Combining Theorem 1.2 with an earlier result of the authors [DKS25, Theorem 1.1], leads to the following corollary proved in Section 4:
@@ 1.1. Outline of contents
The paper is organized as follows: In Section 2 we discuss relations between different, related, properties of skein modules. In Section 3 we prove Theorem 1.3. Section 4 is devoted to the proof of Theorem 1.2. First, using our earlier work [DKS25, Theorem 1.1] and Theorem 1.3, we reduce the proof of Theorem 1.2 to showing that the skein module $\mathcal{S}(M,\mathbb{Z}[A^{\pm 1}])$ is finitely generated for any Seifert fibered space over $S^{2}$ with at most three exceptional fibers and with a non-zero Euler number. We prove the latter statement by skein-theoretic techniques. In Section 5, we compute $|X(M)|$ for all non-Haken Seifert fibered 3-manifolds and, in particular, we prove Theorem 1.5. In the case where $p_{1},p_{2},p_{3}$ are weakly coprime we also compute a basis of $\mathcal{S}(M)$.
@@ 1.2. Acknowledgement
During the course of this work Detcherry was partially supported by the projects ”NAQI-34T” (ANR-23-ERCS-0008) and by the project “CLICQ” of the Région Bourgogne Franche Comté. Kalfagianni was partially supported by the NSF grants DMS-2004155 and DMS-2304033, and Sikora was partially supported by the Simons Foundation grant 957582.
@@ 2. Relations between different properties of skein modules
    $\mathcal{S}(M)$ has no $(A+1)$- nor (A1)(A-1)-torsion over $R=\mathbb{Z}[A^{\pm 1}]$.
@@
The implication $(3)\Rightarrow(6)$ follows from the main theorem of [DKS25]. Indeed, if $\mathcal{S}(M)$ is tame, then $\dim_{Q(A)}(\mathcal{S}(M,\mathbb{Q}(A)))\geq|X(M)|,$ therefore $X(M)$ is finite.
For reduced $\mathcal{X}(M)$, the implication $(3)\Rightarrow(5)$ is the last part of [DKS25, Theorem 3.1].
Finally, if $\mathcal{S}(M,\mathbb{Z}[A^{\pm 1}])$ is generated by $|X(M)|$ elements, then it is isomorphic to $\mathbb{Z}[A^{\pm 1}]^{|X(M)|}/I$ where $I$ is some $\mathbb{Z}[A^{\pm 1}]$-submodule of $\mathbb{Z}[A^{\pm 1}]^{|X(M)|}.$ If $I$ is not the trivial submodule, then we would have that $S(M,\mathbb{Q}(A))=\left(\mathbb{Z}[A^{\pm 1}]^{|X(M)|}/I\right)\underset{\mathbb{Z}[A^{\pm 1}]}{\otimes}\mathbb{Q}(A)$ would have dimension <|X(M)|<|X(M)| over $\mathbb{Q}(A),$ which contradicts [DKS25, Theorem 3.1]. Hence, we must have $I=0,$ and $(7)\Rightarrow(1).$
@@ 3. Haken Seifert fibered manifolds
[…]
@@
The representation $\rho$ now satisfies all relations in the presentation 1, and furthermore, maps each $c_{1}\ldots c_{i}$ for $2\leq i\leq n-2$ to an element of angle $\varphi_{i}.$
We note that Lemma 3.3 actually implies that the dimension of $X(M)$ is at least n3n-3 for $M$ a Seifert manifold over $S^{2}$ with $n\geq 4$ exceptional fibers. We will however not need this fact, but just that $X(M)$ is infinite. The proof of Proposition 3.1 now reduces to the other case, manifolds which fiber over $\mathbb{R}\mathbb{P}^{2}$ with at least $2$ exceptional fibers:
@@
Since $\frac{q_{1}}{p_{1}}$ and $\frac{q_{2}}{p_{2}}$ are not integers, $\sin\left(\frac{q_{1}}{p_{1}}\pi\right)\sin\left(\frac{q_{2}}{p_{2}}\pi\right)\neq 0.$ Consequently, for different choices of $v_{1},v_{2}$, the trace $\mathrm{Tr}(\rho(c_{1}c_{2}))$ can take any value in the interval $\alpha-\beta,\alpha+\beta]$, for
@@
Proof of implication $\Leftarrow:$ Since $X(S^{2}\times S^{1})$ is infinite, we can assume that $M$ is Haken. Let $(M,\pi,B)$ be a Seifert fibered space structure on $M$ where $B$ is the orbifold of the fibration and $\pi:M\longrightarrow B$ is the canonical projection. Recall that $B$ is a closed surface that may be orientable or non-orientable. Since $\pi$ induces a surjection of $\pi_{1}(M)\to\pi_{1}(B)$, if $B\neq S^{2},\mathbb{R}\mathbb{P}^{2}$, then $H_{1}(M)$ projects onto $\mathbb{Z}$, implying an infinite $X(M)$ in this case.
@@ 4. Non-Haken Seifert fibered manifolds
(c) $\Rightarrow$ (a): By [DKS25, Theorem 1.1], we have $|X(M)|\leq\dim_{\mathbb{Q}(A)}\mathcal{S}(M)$ and since by [GJS23], the dimension $\dim_{\mathbb{Q}(A)}\mathcal{S}(M)$ is finite, we conclude that $X(M)$ is finite. Now by Theorem 1.3, $M$ has to be non-Haken, implying it is either reducible or it contains an incompressible surface. ∎
@@ 4.1. The torus skein algebra and the Frohman-Gelca basis
Since multicurves on $T$ without contractible components form a basis of $\mathcal{S}(T,\mathbb{Z}[A^{\pm 1}])$, the skeins $(p,q)_{T}$ for $\pm(p,q)\in\mathbb{Z}^{2}/_{\{\pm 1\}}$ form a basis of $\mathcal{S}(T,\mathbb{Z}[A^{\pm 1}]).$
@@ 4.3. The skein module $\mathcal{S}(S_{0,3}\times S^{1},\mathbb{Z}[A^{\pm 1}]))$
Note that the remaining vectors, $v_{i}$, for i=3,,10i=3,...,10, have their first two components $(k_{1},l_{1}).$ Hence, going from $v_{1}$ to any of them decreases $c_{1}/p_{1}$ by $(p_{1}+q_{1})/p_{1}=1+\varepsilon_{1}.$ Therefore, by (7), it is enough to show that going from $v_{1}$ to one of these vectors increases $c_{2}/p_{2}+c_{3}/p_{3}$ by at most $\max(|\varepsilon_{2}-1|+|\varepsilon_{3}|,|\varepsilon_{2}|+|\varepsilon_{3}+1|)$.
@@
Now we complete the proof of Theorem 4.1 as a corollary of Propositions 4.3 and $4.$
@@
Hence, the above inequality follows from the fact that $|\varepsilon_{3}+1|$ is either $\varepsilon_{3}-1$ or is less than $1$. Recall that $\mathcal{S}(M)$ is generated by the elements $\bar{L}_{v}$ for $v\in\mathbb{Z}^{6}.$ By Propositions 4.3 and 4.4 (which holds by the above inequality), we see that $\mathcal{S}(M)$ is generated by the elements $\bar{L}_{v}$ satisfying
@@
Combining Theorem 1.2 and an earlier result of the authors [DKS25, Theorem 1.1], leads to Corollary 1.4 which we restate here:
@@
It is a direct consequence of Theorem 1.2, combined with an earlier result of the authors [DKS25, Theorem 1.1]. Although Theorem 1.2 implies tameness for irreducible $3$-manifolds only, the above statement does hold the reducible Seifert $3$-manifolds, $S^{2}\times S^{1}$ and $\mathbb{R}\mathbb{P}^{3}\#\mathbb{R}\mathbb{P}^{3}$. In the first case, $\dim_{\mathbb{Q}(A)}\mathcal{S}(S^{2}\times S^{1})=1$ by [HP95] and in the latter,
@@ 5. Character varieties of non-Haken Seifert manifolds
Let $R(M):=\mathrm{Hom}(\pi_{1}(M),\mathrm{SL}_{2}(\mathbb{C}))$. A character $\chi$ of $M$ is called abelian if it is the trace of a diagonal representation $\rho\in R(M)$ and central if $\rho$ takes values in the center $\{\pm I\}$ of $SL(2,\mathbb{C})$. A character is called irreducible if it is the trace of an irreducible $\rho\in R(M)$.
@@ 5.1. Abelian characters
where $\lambda\in\mathrm{Hom}(\pi_{1}(M),\mathbb{C}^{*})\simeq H_{1}(M,\mathbb{Z}).$ Two such representations $\rho_{\lambda}$ and $\rho_{\lambda^{\prime}}$ are conjugated if and only if $\lambda=(\lambda^{\prime})^{\pm 1}$ (as maps $\pi_{1}(M)\rightarrow\mathbb{C}^{*}$). Since one has $\lambda=\lambda^{-1}$ if and only if $\lambda$ takes values in $\{\pm 1\},$ that is, $\lambda\in H_{1}(M,\mathbb{Z}/2\mathbb{Z}),$ the statement follows.
@@
For the second formula, note that for each pair $(p_{i},q_{i})$ at most one of the components is even. So the matrix of above presentation is non-zero mod $2$ and the dimension of $H_{1}(M,\mathbb{Z}/2\mathbb{Z})$ over $\mathbb{Z}/2\mathbb{Z}$ is at most $2.$ The dimension is non-zero if and only if the determinant of this matrix is 0 mod $2.$ For the dimension to be $2$ one needs all of the $2\times 2$ minors: $p_{1}q_{2},\ q_{1}p_{2},\ p_{1}p_{2},\ -q_{1}p_{3}-p_{1}q_{3},\ -q_{1}p_{3},\ -p_{1}p_{3},\ -q_{2}p_{3},\ -q_{2}p_{3}-p_{2}q_{3},\ p_{2}p_{3}$ to vanish in $\mathbb{Z}/2\mathbb{Z}$. The later happens if and only if $p_{1},p_{2}$ and $p_{3}$ are even. ∎
@@
where $m$ is the number of $\varphi\in\mathrm{Hom}(\pi_{1}(M),\mathbb{C}^{*})$ with $\varphi(h)=\pm 1,$ and $m_{i}$ is the number of $\varphi\in\mathrm{Hom}(\pi_{1}(M),\mathbb{C}^{*})$ such that $\varphi(h)=\varphi(c_{i})=\pm 1.$ (Note that if two of $c_{1},c_{2},c_{3}$ are mapped to $\pm 1,$ so is the third one). By Lemma 5.1, $m=|H_{1}(M)/(2h)|$. Eliminating $c_{3}$ from the presentation of $H_{1}(M)/(2h)$ that with respect to the ordered set of generators $\{h,c_{1},c_{2}\}$ the presentation matrix is
@@
If $H_{1}(M,\mathbb{Z})$ is $2$-torsion, then $x_{M}=0$ because all abelian representations have values in $\{\pm I\}$ in this case.
@@ 5.2. Irreducible characters
Each $\chi\in X_{2}(M)$ is either irreducible or exceptional abelian. In either case, $\rho(h)=I$ and $\chi$ is determined by its values on $F_{2}=\langle c_{1},c_{2},c_{3}|c_{1}c_{2}c_{3}=1\rangle$. Since $F_{2}$ is free of rank two, $\chi$ is determined by $(\chi(c_{1}),\chi(c_{2}),\chi(c_{3}))\in\mathbb{C}^{3}$. Thus $T_{+}$ is 11-11. Furthermore, each such value in $\mathbb{C}^{3}$ determines a character of $F_{2}$. In particular, each value in $C_{p_{1}}^{+}\times C_{p_{2}}^{+}\times C_{p_{3}}^{+}$ corresponds to a character of $\langle c_{1},c_{2},c_{3}|c_{1}^{p_{1}}=c_{2}^{p_{2}}=c_{3}^{p_{3}}=c_{1}c_{2}c_{3}=1\rangle$. Thus $T_{+}$ is onto. The same argument shows that $T_{-}$ is a bijection.